# MatCLIP: Light- and Shape-Insensitive Assignment of PBR Material Models

### v1

@@ 6. Conclusion
This paper introduced MatCLIP, a novel framework for robust material classification and assignment to 3D objects. Our method achieves state-of-the-art performance in material classification, leveraging embeddings that are less sensitive to variations in shape and lighting. By training on a diverse dataset of renderings, MatCLIP demonstrates the ability to handle complex material representations, achieving a Top-1 accuracy of 76.69% on MatSynth, significantly outperforming prior top methods such as PhotoShape and MatAtlas. This contribution establishes a new SOTA for robust material classification in diverse geometric and lighting conditions.
Figure 6. Layout of the user study on MTurk. Workers were asked to choose the more plausible material assignment. We compare our method to the procedural assignments from 3DCoMPaT++.
Beyond classification, we demonstrated MatCLIP’s practical application in assigning materials to 3D objects, particularly in scenarios where ground-truth materials are unknown. By integrating Latent Diffusion Models (LDMs) to generate target images, our approach assigns coherent and visually plausible materials, addressing a key challenge in procedural modeling. Qualitative evaluations show that MatCLIP can generalize to unseen material distributions, effectively bridging the gap between synthetic outputs and high-resolution PBR material databases.
By building on publicly available resources like MatSynth and 3DCoMPaT++, MatCLIP ensures accessibility and reproducibility, providing an open framework for the community.

### v2

@@ 6. Conclusion
This paper introduced MatCLIP, a framework for robust material classification and assignment to 3D objects. By training on diverse renderings, MatCLIP learns embeddings invariant to shape and lighting, achieving a Top-1 accuracy of 76.69% on MatSynth and significantly outperforming prior methods.
Beyond classification, MatCLIP supports plausible material assignments to 3D shapes using LDM-generated target images, making a connection between procedural modeling and high-quality PBR materials. By relying on public resources like MatSynth and 3DCoMPaT++, our approach ensures reproducibility and accessibility for the community.
Acknowledgements.
We thank the anonymous reviewers for their constructive comments. The work is supported by funding from KAUST - Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program.